the problem is:
3a-4b=1
12a-b= -11
this is as far as i have gotten...
3a-4b=1
4(12a-b= -11)
________________
3a-4b=1
_ 48a-4b= -44
____________
45a= -43
then i don't know.. i have to solve for a and b
thank you so much for helping me!!!!

Answer by RAY100(1637)

1) 3a -4b = 1
2) 12a - b = -11
mult (2) by 4
2a) 48a -4b = -44
subst (1) from (2a)
45a = -45
a = (-1)
subst in (1)
3(-1) - 4b = 1
-3 -4b =1
-4 = 4b
b= (-1)
check
3 (-1) - 4(-1) =1
1 =1 ok
12(-1) - (-1) = -11
-11 = -11 ok